x+5 3x-1, 4x-2, perimeter=66 find the length of the shortest side
Answer by jim_thompson5910(35256)

Hint:

Add up the three sides and set that sum equal to the perimeter

(Side1) + (Side2) + (Side3) = Perimeter

(x+5) + (3x-1) + (4x-2) = 66

x+5 + 3x-1 + 4x-2 = 66

8x + 2 = 66

I'll let you finish